zayy:

aluminum metal reacts with sufure to produce aluminum sulfide according to this balanced edaution: 2al+3s forms Al2S3. How many atoms of aluminum will react completely with 1.33x10^24 atoms of sulfur

zayy:

@nothnxx wrote:
i think we need 13 atoms of aluminium to react completely with 1.33*10^24atome of sulphur
why would we need 13

Nothnxx:

wait a sec

Nothnxx:

how did you get this much atomes 1.33*10^24 atoms of sulphur?

Nothnxx:

bcs atomic no. of sulphur is 16.

zayy:

@nothnxx wrote:
how did you get this much atomes 1.33*10^24 atoms of sulphur?
i think it was when my teacher balanced everything out

Nothnxx:

@zayy wrote:
@nothnxx wrote:
how did you get this much atomes 1.33*10^24 atoms of sulphur?
i think it was when my teacher balanced everything out
alr

Nothnxx:

wait 2 mins

zayy:

@nothnxx wrote:
wait 2 mins
ok

Nothnxx:

don't you remember how he get 1.33*16^24 huh?

Nothnxx:

bcs there is sum concept of coefficient and all and that i didnt studies yet so

Nothnxx:

studied*

Nothnxx:

i think we gotta do =atomic no. of Sulphur(S) is 16 therefore,how many sulphur they are taking in balancing the qequation so .:16*3=48 .:13*2=26 .:same for aluminium to do take atomic no.then into with 2 as we know how many atoms we need for sulphur (1.33*16^24) so we need to find how many atoms we need for aluminium so i think we gotta aplly this formula:- atomic no. (S)/atomic no. (Al)*1.33*16^24= by solving the equation i think you'll get the answer

Nothnxx:

am i right?

Nothnxx:

and also ur question is wrong it is 1.33 x 1024 atoms of sulfur

zayy:

@nothnxx wrote:
i think we gotta do =atomic no. of Sulphur(S) is 16 therefore,how many sulphur they are taking in balancing the qequation so .:16*3=48 .:13*2=26 .:same for aluminium to do take atomic no.then into with 2 as we know how many atoms we need for sulphur (1.33*16^24) so we need to find how many atoms we need for aluminium so i think we gotta aplly this formula:- atomic no. (S)/atomic no. (Al)*1.33*16^24= by solving the equation i think you'll get the answer
Ohh okkk

Nothnxx:

@zayy wrote:
@nothnxx wrote:
i think we gotta do =atomic no. of Sulphur(S) is 16 therefore,how many sulphur they are taking in balancing the qequation so .:16*3=48 .:13*2=26 .:same for aluminium to do take atomic no.then into with 2 as we know how many atoms we need for sulphur (1.33*16^24) so we need to find how many atoms we need for aluminium so i think we gotta aplly this formula:- atomic no. (S)/atomic no. (Al)*1.33*16^24= by solving the equation i think you'll get the answer
Ohh okkk
well im not sure i given the answer right or not but i kinda thinnk i m right wellask ur teaher how to do then let me know